OpenStudy (anonymous):

A baseball player slides into third base with an initial speed of 7.80 m/s. If the coefficient of kinetic friction between the player and the ground is 0.43, how far does the player slide before coming to rest?

OpenStudy (anonymous):

The player stops sliding when the work done by friction = initial kinetic energy\[1/2mv^2=\mu N * s=\mu mg * s\\s=\frac{v^2 }{ 2\mu g }\]For g = 9.8m/s^2, he slides 7.2 m

OpenStudy (anonymous):

(Technically the work done by friction is negative, so I should have said the *magnitude* of the work done by friction)

OpenStudy (anonymous):

Okay thank you! this helped a lot